### v2

@@
3 Tereshchenkivs’ka Street, Kyiv-4, Ukraine, 01024, and
Universit´a del Piemonte Orientale,
@@ 1 Introduction
Second order integrals of motion for 2d PDM SEs are perfectly classified [20, 21, 22, 23]. The majority of such systems admits also at least one continuous Lie symmetry. The two dimensional second-order (maximally) superintegrable systems for Euclidean 2-space had been classified even algebraic geometrically [24].
@@
Surely, the maximally superintegrable systems are very important and interesting. In particular, they admit solutions in multi separated coordinates [31, 32, 33, 34]. On the other hand, there are no reasons to ignore the PDM systems which admit second order integrals of motion but are not necessary maximally superintegrable. And just such systems are the subject of our study.
@@
It was shown in [17] the PDM Schrödinger equation can admit six, four, three, two or one parametric Lie symmetry groups. In addition, there are also such equations which have no Lie symmetry. In other words, there are six well defined classes of such equations which admit n-parametric Lie groups with $n=6,4,3,2,1$ or do not have any Lie symmetry. And it is a natural idea to search for second order integrals of motion consequently for all these classes.
@@ 2 Formulation of the problem
$H=p_{a}f({\bf x})p_{a}+V({\bf x}),$ (2)
${\bf x}=(x^{1},x^{2},x^{3}),$ $p_{a}=-i\partial_{a}$, $V({\bf x})$ and $f({\bf x})=\frac{1}{2m({\bf x})}$ are functions associated with the potential and inverse PDM, and summation from 1 to 3 is imposed over the repeating index $a$.
@@
$\mu^{ab}V_{b}-f\eta_{a}=0$ (11)
where $\delta^{bc}$ is the Kronecker delta, $f_{n}=\frac{\partial f}{\partial x_{n}},\ \mu^{an}_{n}=\frac{\partial\mu^{an}}{\partial x_{n}}$, etc., and summation is imposed over the repeating indices $n$ over the values $n=1,2,3$.
Equations (9) , (10) and (11) give the necessary and sufficient conditions for commutativity of operators $H$ (2) and $Q$ (7) [39].
@@ 3 Evolution of the determining equations
$\left(\mu^{nn}_{a}+2\mu^{na}_{n}\right)\tilde{M}+5\mu^{an}\tilde{M}_{n}+R_{a}=0$ (16)
@@
$f=\frac{2}{M},\ V=\frac{\tilde{M}}{M}$ (17)
@@ 6.1 Scalar integrals of motion
Since the searched integrals of motion include bilinear combinations of the generators $S_{\mu\nu}$ of the extended Lorentz group SO(1,4), it is relatively easy to select such of them which commute with $S_{12}=L_{3}$ and so are scalars with respect to the rotations. First, all $S_{\mu\nu}$ with $\mu,\nu\neq$ 1 and 2 commute with $S_{12}$ and so their bilinear combinations are scalars. There exactly three such $S_{\mu\nu}$, namely, $S_{04},S_{03}$ and $S_{43}$. Surely, this list can be added by $S_{12}$. Thus we have indicated the following scalars:
$\begin{array}[]{c}S_{04}^{2},\ S_{03}^{2},\ S_{34}^{2},\ S_{12}S_{04},\ S_{12}S_{03},\ S_{12}S_{34},\\ S_{34}S_{04}+S_{04}S_{34},\ S_{03}S_{04}+S_{04}S_{03},\ S_{43}S_{03}+S_{03}S_{43}.\end{array}$ (45)
The above presented list does not include the squared generator $S_{12}$ which commutes with all the considered Hamiltonians by definition.
One more collection of scalars which includes the sums w.r.t. the repeating indices $n=1,2$ and $m=1,2$ is presented in the following formulae:
$\begin{split}&S_{na}S_{nb}+S_{nb}S_{na}\end{split}$ (46)
and
$\begin{split}&\varepsilon_{3nm}(S_{na}S_{mb}+S_{mb}S_{na})\end{split}$ (47)
where $a,b=0,3,4$.
Equation (1) including arbitrary parameters (3) is transparently invariant with respect the inversion of variables $x_{1}$ or $x_{2}$. The consequence of this observation is that its scalar integrals of motion can be decoupled to two linearly independent parts including proper scalars and pseudo scalars. The proper scalars are collected in (45) and (46) while the pseudo scalars are present in (47).
The generic scalar integral of motion is a linear combination of operators (45), (46), and the term $P_{a}g(\tilde{r},x_{3})P_{a}$ (refer to (28)). The pseudo scalar ones are the linear spans of operators (47) added by the same term. Moreover, many terms of this linear combination should be omitted in view of relations (23) and (29)-(32). As a result we will obtain the reduced set of scalars which we write in terms of operators (21):
$P_{3}L_{3},\ DL_{3},\{L_{3},K_{3}\},$ (48)
$P_{3}^{2},\{P_{3},K_{3}\},\ D^{2},\ K_{3}^{2},\ \{P_{3},D\},\ \{K_{3},D\}.$ (49)
Notice that equation (46) collects the pseudo scalars while the proper scalars are presented in (49).
One more way to specify the scalar integrals of motion is to use the formal expressions for Killing tensors given by relations (12)-(41). Such integrals are generated by the mentioned tensors in two cases: when the only nonzero parameters $\lambda^{a}_{n}$ and $\lambda^{ab}_{n}$ correspond to $a=3$ and $a=b=3$. One more possibility is $\lambda^{11}=\lambda^{22}\neq 0$, but it is reduced to the previous one in view to the traceless condition for constant tensors $\lambda^{ab}_{n}$. Going over the mentioned values of arbitrary parameters we recover relations (48) and (49).
@@
$\hat{Q}_{2}=(K_{3}+P_{3})L_{3}+P_{a}\tilde{q}_{2}(\tilde{r},x_{3})P_{a},\ \mu^{ab}=\mu_{3}^{ab}+\mu_{8}^{ab},$ (54)
$\hat{Q}_{3}=(K_{3}-P_{3})L_{3}+P_{a}\tilde{q}_{3}(\tilde{r},x_{3})P_{a},\ \mu^{ab}=\mu_{3}^{ab}-\mu_{8}^{ab}$ (55)
@@
Operator (56) is equivalent to (54) and can be ignored in the analysis of integrable PDM systems. However, we cannot ignore it in the case of superintegrable systems.
@@
The next task is to specify all inequivalent proper scalars. In accordance with (48 ) and (51 they belong to the enveloping algebra of algebra so(1,2). Moreover, it follows from the first of equations (30) that the related Casimir operator of algebra so(1,2) takes the following form:
@@
The l.h.s. of relation (60) includes the symmetry operator $L_{3}^{2}$ which commutes with Hamiltonian (2) by definition and the term $P_{a}\tilde{r}^{2}P_{a}$ which can be included to the last term of the generic integral of motion (28). It means that operators $Q_{1},Q_{2}$ and $Q_{6}$ can be treated as linearly dependent and so one of the coefficients $c_{1},c_{2}$ and $c_{3}$ can be nullified without loss of generality.
The next step in the simplification of expression (58) can be made using the equivalence transformations generated by the Lie group whose Lie algebra is a linear span of base elements (50). The transformations generated by $S_{04}$ are the Lorentz transformations which look as:
$\begin{split}&Q_{2}\to Q_{2}\cosh(2\lambda)+Q_{3}\sinh(2\lambda),\\ &Q_{3}\to Q_{3}\cosh(2\lambda)+Q_{2}\sinh(2\lambda),\end{split}$ (61)
$\begin{split}&Q_{4}\to Q_{4}\cosh(\lambda)+Q_{5}\sinh(\lambda)\\ &Q_{5}\to Q_{4}\cosh(\lambda)+Q_{5}\sinh(\lambda).\end{split}$ (62)
In addition, $Q_{6}$ and $Q_{2}$ are transformed to their linear combinations, but such transformations can be ignored in view of relation (60).
Let us use transformations (61) to specify the inequivalent versions of coefficients $c_{2}$ and $c_{3}$ present in formula (58). For $c_{2}^{2}>c_{3}^{2}$ or $c_{2}^{2}<c_{3}^{2}$ we can nullify $c_{3}$ or $c_{2}$ correspondingly, while for the special case $c_{4}=\pm c_{5}$ we can restrict ourselves to the case $c_{4}=-c_{5}$ up to the discrete equivalence transformations (22). So the inequivalent versions of the pairs $(c_{4},c_{5})$ are:
$c_{2}\neq 0,c_{3}=0;\ c_{2}=0,c_{3}\neq 0;\ c_{2}=-c_{3}.$ (63)
Whenever the last version presented in (63) is true we can use transformations (62) to specify the inequivalent versions of constants $c_{4}$ and $c_{5}$ which are analogous to the ones presented in (63):
$c_{4}\neq 0,c_{5}=0;\ c_{4}=0,c_{5}\neq 0;\ c_{4}=-c_{5}$ (64)
provided $c_{2}=-c_{3}.$
The next step is to use the rotation like transformation generated by $S_{43}$ which look as follows:
$\begin{split}&Q_{3}\to Q_{3}\cos(\omega)-Q_{5}\sin(\omega),\\ &Q_{5}\to Q_{5}\cos(\omega)+Q_{3}\sin(\omega),\\ \end{split}$ (65)
$\begin{split}&Q_{6}\to Q_{6}\cos(2\omega)-Q_{4}\sin(2\omega),\\ &Q_{4}\to Q_{4}\cos(2\omega)+Q_{6}\sin(2\omega)\end{split}$ (66)
Let at least one of the constants $c_{3},c_{3}$ is nontrivial than transformations (65) can be used to transform them to the following form:
$c_{5}=0\ \ c_{3}\neq 0.$ (67)
Alternatively, both $c_{5}$ and $c_{3}$ can be equal to zero.
Summarising the results presented in (63), (64) and (67) and taking into account that in view of relation (60) coefficients $c_{1},c_{2},c_{3}$ can be treated as linearly dependent we come to the following inequivalent versions of operators (58) which we present together with the related Killing tensors:
@@ 6.2 Pseudo scalar integrals of motion
Let us start which the pseudo scalar integrals of motion fixed in (53). The related determining equations (14) are reduced to the following form:
@@
and are solved by the following functions:
$M=\frac{\nu x_{3}+F(\tilde{r})}{\tilde{r}^{2}},\ N=\frac{\nu}{2}\varphi$ (77)
where $\varphi=\arctan\left(\frac{x_{2}}{x_{1}}\right)$ is the Euler angle, $\nu$ is the integration constant, and $F(\tilde{r})$ is an arbitrary function of $\tilde{r}=\sqrt{x_{1}^{2}+x_{2}^{2}}.$ The corresponding functions $\tilde{M}$ and $\tilde{N}$ which generate potential $V$ and function $\eta$ are solutions of the same equation and so have the same generic forms as given in (77), but in general with different parameter $\nu$ and different function $F(\tilde{r})$. Substituting the obtained results into equations (18) we come to functions $f$ and $V$ presented in Item 1 of Table 1.
The next pseudoscalar operator which we consider is fixed by relations (54) and (55). The corresponding determining equations (14) for $M=M(\tilde{r}^{2},x_{3})$ and $N=N(\tilde{r}^{2},x_{3})$ are reduced to the form:
@@
where $\partial_{3}=\frac{\partial}{\partial_{x_{3}}},\partial_{4}=\frac{\partial}{\partial_{x_{4}}},x_{4}=\tilde{r}^{2}$.
In accordance with (78) function $N$ depends on $\varphi$ only. Since $M$ by definition does not depend on this variable, to solve equation (79) $N$ should be a linear function, i.e., $N=c\varphi.$ Then equation (79) is reduced to the following one:
@@
Equation (80) is easily integrable. Its solutions are presented in Items 2 and 3 of Table 1. Notice that these solutions are qualitative different for different sighs before 1 in the formulae presented above. And these signs are the same as signs for $P_{3}$ in equations (54) and (55).
@@
where $F(\theta)$ and $G(\theta)$ are arbitrary functions, $\varphi$ and $\theta$ are the Euler angle, $\mu$ and $\nu$ are arbitrary parameters.
@@ 6.3 True scalar integrals of motion
Two more particular solutions are:
$M=\frac{1}{\tilde{r}^{2}}\ \text{ and }\ M=\frac{x_{3}^{2}+1}{\tilde{r}^{4}}.$ (125)
Notice that linear combinations of solutions (123), (124) and (125) also solve equations (14), see Item 10 of Table 1.
The last class of symmetries which we are supposed to study is represented in equation (74). The corresponding determining equations (105) are the most complicated, but in many aspects they are analogous to ones requested for the systems admitting integrals of motion of type (73).
Excluding $N$ we come to the following compatibility condition for the system (105):
$\begin{array}[]{l}((x_{4}-x_{3}^{2}+n)^{2}-n^{2}\pm 1-4x_{3}^{2}x_{4})\partial_{3}\partial_{4}-(n+x_{4}-x_{3}^{2})(x_{3}\partial_{3}\partial_{3}-4x_{3}x_{4}\partial_{4}\partial_{4})\\ +3(n+x_{4}-3x_{3}^{2}r)\partial_{3}-(8n-8x_{3}^{2}+20x_{4})x_{3}\partial_{4}-12x_{3})M=0.\end{array}$ (128)
This frighteningly looking equation appears to be equivalent to the d’Alambert one if we choose the new independent and dependent variables
[…]
As a result equation (128) takes the canonical form
$\partial_{y}\partial_{\tilde{y}}\hat{M}=0$ (129)
and so the related mass function has the following form:
$M=\frac{F(y)+G(\tilde{y})}{x_{4}\sqrt{z_{\pm}^{2}+2x_{3}^{2}x_{4}}}$
with arbitrary functions $F(.)$ and $G(.)$.
For any particular functions $F(.)$ and $F_{2}(.)$ we can find the corresponding functions $N$ by direct integration of equations (105). Moreover, for some rather extended classes of the arbitrary functions it is possible to present the related functions $N$ in closed form like it was done in equations (122)-(125). In particular we can set:
$M=M_{m}=\frac{1}{x_{4}\sqrt{z_{\pm}^{2}+2x_{3}^{2}x_{4}}}(y^{m}+(-1)^{m+1}\tilde{y}^{n})$ (130)
where $n$ are natural numbers. Then the corresponding functions $N=N_{m}$ can be found in the closed form for arbitrary $m$ by directs integration of equations (105):
$N_{m}=2^{m-1}x_{3}^{2}M_{m-1}.$ (131)
We again have a countable set of exact solutions whose linear combinations $\sum_{m}c_{m}M_{m}$ are solutions also. The corresponding potentials looks as follows:
$V=\frac{\sum_{m}\tilde{c}_{m}(y^{m}+(-1)^{m+1}\tilde{y}^{m})}{\sum_{m}c_{m}(y^{m}+(-1)^{m+1}\tilde{y}^{m})}.$ (132)
The obtained in this way solutions are represented in Items 11 and 12 of Table 1. Solutions corresponding to $n^{2}-1<0$ which are represented in Items 13 and 14 of the same table can be obtained in analogous way.
@@
In Table 1 and the following Tables 2, 3, 4 $F(.),G(.),\tilde{F}(.)$ and $\tilde{G}(.)$ are arbitrary functions, $g(F,G)$ and $\eta(\tilde{F},\tilde{G})$ are solutions of equation (101), (105) with given $F_{1}(x),F_{2}(\tilde{x})$ and $G_{1},G_{2}$, $c_{1},c_{2},..$ are arbitrary real parameters. In addition we use the following notations which also will be used in all tables:
@@
The mysterious term $\hat{\eta}$ appears as a result of the following identity:
$K_{3}^{2}=p_{a}\mu_{8}^{ab}p_{b}+\hat{\eta}$
where $\mu_{9}^{ab}$ is the Killing tensor (41) with $\lambda_{9}^{3}\neq 0$ , and it is the cost paid for our desire to represent the integrals of motion via the anticommutators of operators (21).

### v3

@@
3 Tereshchenkivs’ka Street, Kyiv-4, Ukraine, 01024,
and Universit´a del Piemonte Orientale,
@@ 1 Introduction
Second order integrals of motion for 2d PDM SEs are perfectly classified [20, 21]. The majority of such systems admits also at least one continuous Lie symmetry. The two dimensional second-order (maximally) superintegrable systems for Euclidean 2-space had been classified even algebraic geometrically [22].
@@
Surely, the maximally superintegrable systems are very important and interesting. In particular, they admit solutions in multi separated coordinates [29, 30, 31]. On the other hand, there are no reasons to ignore the PDM systems which admit second order integrals of motion but are not necessary maximally superintegrable. And just such systems are the subject of our study.
@@
It was shown in [17] the PDM Schr¨odinger equation can admit six, four, three, two or one parametric Lie symmetry groups. In addition, there are also such equations which have no Lie symmetry. In other words, there are six well defined classes of such equations which admit n-parametric Lie groups with $n=6,4,3,2,1$ or do not have any Lie symmetry. And it is a natural idea to search for second order integrals of motion consequently for all these classes.
@@ 2 Formulation of the problem
$H=\frac{1}{4}(m^{\alpha}p_{a}m^{\beta}p_{a}m^{\gamma}+m^{\gamma}p_{a}m^{\beta}p_{a}m^{\alpha})+\hat{V}(x)$ (2)
where $p_{a}=-i\partial_{a}$, $m=m({\bf x})$ is a function of spatial variables ${\bf x}=(x_{1},x_{2},x_{3})$, associated with the position dependent mass, $\alpha,\beta$ and $\gamma$ are the so called ambiguity parameters satisfying the condition $\alpha+\beta+\gamma=-1$, and and summation from 1 to 3 is imposed over the repeating index $a$.
There are various physical speculations how to fix the ambiguity parameters in the particular models based on equation (2) . However, the systems with different values of these parameters are mathematically equivalent up to redefinition of potentials $V({\bf x})$. To simplify the following calculations we will fix them in the following manner: $\beta=0,\gamma=\alpha=-\frac{1}{2}$ and denote $f=f({\bf x})=\frac{1}{2m({\bf x})}$. As a result hamiltonian (2) is reduced to the following form
$H=f^{\frac{1}{2}}p_{a}p_{a}f^{\frac{1}{2}}+V({\bf x}).$ (3)
In addition we consider the version $\alpha=\gamma=0,\beta=-1$ which corresponds to the following form of hamiltonian (2):
$H=p_{a}fp_{a}+\hat{V}.$ (4)
Operators (4) and (3) are equal one to another provided potentials $V=V({\bf x})$ and $\hat{V}=\hat{V}({\bf x})$ satisfy the following condition:
$V=\hat{V}+V^{k}$ (5)
where
$V^{(k)}=\frac{1}{4f}((\partial_{1}f)^{2}+(\partial_{2}f)^{2}+(\partial_{3}f)^{2})-\frac{1}{2}\Delta f$ (6)
and $\Delta$ is the Laplace operator.
Formula (6) represents an example of kinematical potentials which can be reduced to zero by the rearranging the ambiguity parameters.
@@
$\mu^{ab}V_{b}-f\eta_{a}+F^{a}=0$ (15)
where $\delta^{bc}$ is the Kronecker delta, $f_{n}=\frac{\partial f}{\partial x_{n}},\ \mu^{an}_{n}=\frac{\partial\mu^{an}}{\partial x_{n}}$, etc., and summation is imposed over the repeating indices $n$ over the values $n=1,2,3$. Moreover,
$F^{a}=(f\mu^{ak}_{km}-\mu^{mn}f_{na})_{m}$ (16)
for the Hamiltonian of form (4) and
$F^{a}=(f\mu^{ak}_{km}-\mu^{mn}f_{na})_{m}+\mu^{ab}V^{(k)}_{b}$ (17)
for the Hamiltonian of form (5).
Equations (13) , (14) and (15) give the necessary and sufficient conditions for commutativity of operators $H$ (3) and (4) with $Q$ (11) [34].
@@ 3 Evolution of the determining equations
$\left(\mu^{nn}_{a}+2\mu^{na}_{n}\right)\tilde{M}+5(\mu^{an}\tilde{M}_{n}+R_{a})=0$ (22)
@@
$f=\frac{1}{M},\ V=\frac{\tilde{M}}{M}$ (23)
@@ 6.1 Scalar integrals of motion
The scalar integrals of motion are generated by the Killing tensors (39)-(47) with the following nontrivial parameters $\lambda^{a}_{n}$ and $\lambda_{m}^{ab}$:
$\lambda_{n}^{3}\neq 0,\ n=2,4,5,7;\ \lambda_{m}^{33}\neq 0,\ m=1,3,6,8,9$ (49)
while the remaining parameters should be identically zeros. These scalars can be expressed via bilinear combinations of generators (27) in the following manner:
$S_{3}=P_{3}L_{3}+...,\ S_{4}=DL_{3}+...,S_{8}=\{L_{3},K_{3}\}+...,$ (50)
$\begin{split}&S_{1}=P_{3}^{2}+...,S_{2}=\{P_{3},D\}+...,S_{6}=\{P_{3},K_{3}\}+...,\\ &S_{5}=D^{2}+...,\ S_{7}:=\{K_{3},D\}+...,\ S_{9}:=K_{3}^{2}+...\end{split}$ (51)
where the dots denote the differential operators $S_{0}=P_{a}g_{n}(\tilde{r},x_{3})P_{a}$ generated by $\mu_{0}$ (18) and functions $\eta_{n}$ present in definition (11), and the numbers presented as subindices coincide with the number of the Killing tensor in (39-(47) which generate the scalar.
All operators (50) and (51) are invariant with respect to rotations around the third coordinate axis. However, they have different transformation properties with respect to reflections (32) and (33). Namely, operators (50) change their sign and so are pseudo scalars while operators (51) remain unchanged and so are true or proper scalars.
The generic scalar integral of motion is a linear combinations of operators (50) and (51). Surely, such a linear combination cannot include both true and pseudo scalars, and we have one more separation of our integrals to two subclasses.
@@
$\hat{Q}_{2}=(K_{3}+P_{3})L_{3}+P_{a}\tilde{q}_{2}(\tilde{r},x_{3})P_{a}+\eta_{2},\ \mu^{ab}=\mu_{3}^{ab}+\mu_{8}^{ab},$ (56)
$\hat{Q}_{3}=(K_{3}-P_{3})L_{3}+P_{a}\tilde{q}_{3}(\tilde{r},x_{3})P_{a}+\eta_{3},\ \mu^{ab}=\mu_{3}^{ab}-\mu_{8}^{ab}$ (57)
@@
Operator (58) is equivalent to (56) and can be ignored in the analysis of integrable PDM systems. However, we cannot ignore it in the case of superintegrable systems when it can appear in a combinations with other symmetries.
@@
The next task is to specify all inequivalent proper scalars. In accordance with (50 ) and (53) they belong to the enveloping algebra of algebra so(1,2). Moreover, it follows from the first of equations (36) that the related Casimir operator of algebra so(1,2) takes the following form:
@@
The l.h.s. of relation (62) includes the symmetry operator $L_{3}^{2}$ which commutes with Hamiltonian (3) by definition and the term $P_{a}\tilde{r}^{2}P_{a}$ which can be included to the last term of the generic integral of motion (34). It means that operators $Q_{1},Q_{2}$ and $Q_{6}$ can be treated as linearly dependent and so one of the coefficients $c_{1},c_{2}$ and $c_{3}$ can be nullified without loss of generality. Thus the generic linear combination (61) includes five terms only.
Rather tedious speculations with using the mentioned in the above equivalence relations and identities (35) (see the preprint [37]) for the details) make it possible to specify the following inequivalent scalar symmetries (61) which we present together with the related Killing tensors:
@@ 6.2 Pseudo scalar integrals of motion
First we present the related determining equations which take the following form:
@@
for $\hat{Q}_{1},$
@@
for $\hat{Q}_{2}$ and $\hat{Q}_{3}$ were $\hat{Q}_{1}-\hat{Q}_{3}$ are operators (55)- (57), $\partial_{3}=\frac{\partial}{\partial_{x_{3}}},\partial_{4}=\frac{\partial}{\partial_{x_{4}}},x_{4}=\tilde{r}^{2}$.
Equations (71) are easily integrable and are solved by the following functions:
$M=\frac{\nu x_{3}+F(\tilde{r})}{\tilde{r}^{2}},\ N=\frac{\nu}{2}\varphi\ \text{ for }\hat{Q}_{1},$ (74)
where $\varphi=\arctan\left(\frac{x_{2}}{x_{1}}\right)$ is the Euler angle, $\nu$ is the integration constant, and $F(\tilde{r})$ is an arbitrary function of $\tilde{r}=\sqrt{x_{1}^{2}+x_{2}^{2}}.$ The corresponding functions $\tilde{M}$ and $\tilde{N}$ which generate potential $V$ and function $\eta$ are solutions of the same equation and so have the same generic forms as given in (74), but in general with different parameter $\nu$ and different function $F(\tilde{r})$. Substituting the obtained results into (24) we come to functions $f$ and $V$ presented in Item 1 of Table 1.
Equations (72) and (73) are easily solvable also. In accordance with (72) function $N$ depends on $\varphi$ only. Since $M$ by definition does not depend on this variable, to solve equation (73) $N$ should be a linear function, i.e., $N=c\varphi.$ Then equation (73) is reduced to the following one:
@@
Its solutions are presented in Items 2 and 3 of Table 1.
Notice that these solutions are qualitative different for different sighs before 1 in the formulae presented above. And these signs are the same as signs for $P_{3}$ in equations (56) and (57).
@@
where $F(\theta)$ is arbitrary function, $\varphi$ and $\theta$ are the Euler angles, $\mu$ and $\nu$ are arbitrary parameters.
@@ 6.3 True scalar integrals of motion
Notice that linear combinations of solutions (118), (119) also solve equations (20), see Item 10 of Table 1.
The last class of symmetries which we are supposed to study is represented in equation (69). The corresponding determining equations (100) are the most complicated, but in many aspects they are analogous to ones requested for the systems admitting integrals of motion of type (68). The related solutions are represented in Items 11-12 of Table 1, see preprint [37] for more detailed calculations. Solutions corresponding to $n^{2}-1<0$ which are represented in Items 13 and 14 of the same table can be obtained in analogous way.
@@
In Table 1 and the following Tables 2-5 $F(.),G(.),\tilde{F}(.)$ and $\tilde{G}(.)$ are arbitrary functions, $g(F,G)$ and $\eta(\tilde{F},\tilde{G})$ are solutions of equation (96), (100) with given $F_{1}(x),F_{2}(\tilde{x})$ and $G_{1},G_{2}$, $c_{1},c_{2},..$ are arbitrary real parameters.
@@
In addition we use the following notations which also will be used in all tables:
@@
The mysterious term $\hat{\eta}$ appears as a result of the following identity: $K_{3}^{2}=P_{a}\mu_{8}^{ab}P_{b}+\hat{\eta}$ where $\mu_{9}^{ab}$ is the Killing tensor (47) with $\lambda_{9}^{3}\neq 0$ , and it is the cost paid for our desire to represent the integrals of motion via the anticommutators of operators (27).